OpenStudy (anonymous):

1.Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make?

OpenStudy (♪chibiterasu):

Lol my sis did this recently, can we see their work?

OpenStudy (anonymous):

wut uu mean?

OpenStudy (♪chibiterasu):

We need to see the work they did in order to solve.

OpenStudy (anonymous):

2x + 4 = –3x + 14 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 2x + 4 – 2x = –3x + 14 – 2x 5x + 4 = 14 4 = 14 – 5x 5x + 4 + 4 = 14 + 4 4 – 14 = 14 – 5x – 14 5x = 18

OpenStudy (anonymous):

–12 = –5x –12 + 5 = –5x + 5 x = –7 = x

OpenStudy (♪chibiterasu):

The 4th row of Splott's calculation is unnecessary because he adds 4 instead of subtracting which leads to error on the next row. In Fizzle's 6th row there is addition of 5 instead of 5x which leads to an incorrect final row.

OpenStudy (♪chibiterasu):

I have to go to bed soon so I won't have time for explanations, sorry xc